# Analyzing the numerical correctness of branch-and-bound decisions for mixed-integer programming

### v1

@@ 2 Numerical Errors in LP-based Branch and Bound
where we assume rational input data $A\in\mathbb{R}^{m\times n}$, $c\in\mathbb{R}^{n}$, $u,l\in(\mathbb{R}\cup\{\pm\infty\})^{n}$, and $b\in\mathbb{R}^{m}$. The index set $\mathcal{I}\subseteq\{1,\dots,n\}$ defines which decision variables are required to be integer.
MIPs are typically solved by variants of LP-based branch and bound, which dates back to [16] and still forms the backbone of today’s most competitive MIP solvers.
In Section 2.1, we describe briefly the procedure of an LP-based branch-and-bound algorithm. For a more detailed description please refer to [1]. In Section 2.3, we give an overview of the techniques we use to obtain an exact evaluation of a node that was solved in floating-point arithmetic. In Section 2.2, we categorize how numerics can impact the critical branch-and-bound decisions in a floating-point solver and how incorrect decisions affect the overall result.
@@ 2.1 Basics of the LP-based Branch-and-Bound Algorithm
The first step in an LP-based branch-and-bound algorithm is to relax the by dropping all integrality constraints. The resulting LP relaxation can be solved by an LP solver. If the LP solution is integer feasible, i.e., all integer variables have an integer solution value, the problem is solved. If the LP relaxation is detected to be infeasible, then also the original MIP is infeasible. Otherwise, the algorithm branches: It chooses an integer variable $x_{i}$ with a fractional value $\hat{x}_{i}$ in the solution of the relaxed problem to create two new sub-problems by taking the original problem and adding the constraint $x_{i}\geq\lceil\hat{x}_{i}\rceil$ and $x_{i}\leq\lfloor\hat{x}_{i}\rfloor$, respectively. This ensures that the optimal solution is still preserved in at least one of the two sub-problems and the current LP solution is no longer feasible in both sub-problems. This process of LP solving and branching is iterated recursively.
If during this process an integer feasible solution is found its objective value can be used to the primal bound, which is the objective value of the best known solution encountered so far. This primal bound can be used to prune, i.e., remove from further consideration all open sub-problems with a greater or equal dual bound, which is given by the objective value of the sub-problem’s LP relaxation.
This recursive procedures produces a search tree with sub-problems as nodes. A node is called leaf if it is not further branched on because it is a) integer feasible, b) LP infeasible, or c) it was pruned. . Figure 1 presents an example of a branch-and-bound tree for a minimization problem. The solution of the relaxed LP at the node is displayed slightly above and to the right of the node, while its objective value is positioned on the left side. The leaf $N_{4}$ is integer feasible and appears to be also optimal. The leaf $N_{3}$ is infeasible and the node $N_{1}$ is cut off since branching on $N_{1}$ does not improve the solution of $N_{4}$ anymore.
@@
branch and bound, where level=1 set branch labels=x_2≤x_2≥, if level=2 set branch labels=x_3≤x_3≥, , [ $1.5$:root:{0;$\frac{1}{2}$;1} [ $2.5$:N_1:{1;0;$\frac{1}{2}$}:0] [ $1.5$:N_2:{0;1;$\frac{1}{2}$}:1 [ :N_3:infeas:0] [2:N_4:{0;1;1}:1] ] ]
@@ 2.2 Types of Numerical Errors in LP-based Branch and Bound
By design, a floating-point solver may produce approximate solutions with slight violations of integrality and constraint feasibility. A simple post-processing step is to first round the value of all integer variables to the nearest integer and second, if continuous variables are present, compute rational values for these such that all constraints are satisfied exactly. If this is certifiably not possible, i.e., if the floating-point solver accepted an assignment for the integer variables as feasible that cannot be completed to an exactly feasible solution, we refer to this as a solution error.
@@
If a floating-point solver uses a lower bound to prune a node, this decision can often be verified a posteriori by computing a safe dual bound that is valid in exact arithmetic, by techniques described in Section 2.3. If this safe bound is greater than or equal to the current primal bound, then the decision to prune the node was correct. Otherwise, if the exact objective value of the LP relaxation is below the primal bound, we label this decision as a bound error. Again, we distinguish two cases. If later during the solving process an exactly feasible solution is found that improves the primal bound and in hindsight justifies pruning the leaf, we call the error a weak bound error, otherwise a strong bound error.
@@
For the last type of error, consider the situation that a floating-point solver declares a node to be integer feasible and produces an approximate solution that can be converted to an exactly feasible solution after rounding the integer assignment. Although we record no solution error here, the decision to terminate the search below this node may still be incorrect if the objective value of the converted, exact solution does not match the dual bound derived by solving the LP relaxation exactly. We refer to this error as a gap error, and as above we distinguish between weak gap errors that are justified in hindsight and strong gap errors.
@@
If the floating-point solver declares a leaf infeasible but there exists a solution to the LP relaxation that is exactly feasible we refer to this situation as an infeasibility error.
@@
Also in the presence of strong bound or strong gap errors, the solution or at least its objective value $z^{*}$ returned by the floating-point solver may still be optimal. Because our analysis produces a safe dual bound at all occurrences of such errors, we can then take the minimum of all these dual bounds and derive a global lower bound $\hat{z}<z^{*}$ for the true optimal objective value. This certifies that the true optimal objective value must lie in the interval $[\hat{z},z^{*}]$. If infeasibility errors occur, i.e., if a leaf is removed although the LP relaxation is feasible in exact arithmetic, we can use the exact objective value of the LP relaxation as a valid dual bound.
@@ 2.3 Checking Numerical Correctness of Floating-point Decisions
A straightforward way to evaluate whether and which of the errors described above occurs at a leaf node is to solve the LP relaxation once again with an exact LP solver and compare the results. For most but small instances, this approach is prohibitively slow. Hence, we use a hierarchy of methods that may be able to verify or refute the correctness of leaf decisions faster.
If the plain floating-point verification fails we apply the following steps:
@@
In [17], Neumaier and Shcherbina introduce safe bounding: We assume we have an approximated dual solution $\{\hat{y},\hat{r_{-}},\hat{r_{+}}\}$ with the error
$\varepsilon=c-A^{T}\hat{y}+I\hat{r}^{+}+I\hat{r}^{-}\;.$ (1)
We replace
$r^{-}_{i}:=\hat{r}^{-}_{i}+\min\{\varepsilon_{i},0\}\textnormal{and }r^{+}_{i}=\hat{r}^{+}_{i}+\max\{\varepsilon_{i},0\}$ (2)
for all variables $i\in N$ and obtain a feasible dual solution $\{\hat{y},r^{-},r^{+}\}$ with $b^{T}\hat{y}-l^{T}r^{+}+u^{T}r^{-}$ as lower bound on the primal objective. This approach requires finite variables bounds $l,u$. Ideally, the variable bounds are small since they impact the quality of the derived lower bound by construction. Safe-bounding can be applied analogously for validating the Farkas proof. This approach is also implemented for validating floating-point LP solutions and explained in more detail in [21].
Rational reconstruction
We apply continued fractions approximation computed by the extended Euclidean algorithm to re-construct an approximated solution to an exact solution. This method is applied to both, the primal and the dual solution, as well as to the Farkas proof. We rely on the implementation of SoPlex [13, 14]. This method has been introduced and improved by [23, 25, 18, 20].
Factorization
If available, we perform a rational LU factorization on the simplex basis achieved from the floating-point LP solver. For this, we use the implementation in SoPlex [13, 14]. This step is only skipped for infeasible nodes since no basis can be obtained from the LP solver.
@@
The last and most time-expensive option to derive the most accurate lower bound on the objective is to call a round-off-error-free LP solver. The open-source solver SoPlex provides a configuration to perform such round-off-error-free LP solver [13, 14]. We refer to this configuration as Exact-SoPlex. This is performed as the last option for all cases.
For obtaining a primal solution from an inaccurate floating-point solution, we also can fix the integer variables to the values of the floating-point solution and solve the LP over the continuous variables in rational arithmetic with Exact-SoPlex. For each primal solution value, it is ensured to lie within its lower and upper bounds.
@@ 3 Computational Study
Guiding question for the design of our experiments were the following: Assuming the final result produced by a floating-point branch-and-bound solver is correct, can we also verify that the reasoning of the individual decisions of the solver is correct? If not, how many of the critical decision would require re-evaluation, and which types of the errors defined in Section 2.2 occur at which frequency? Furthermore, we are interested in the effectiveness of the different post-processing techniques outlined in Section 2.3. Before evaluating our experiments in Section 3.2, we first describe the setup and specify the software used in Section 3.1.
@@ 3.1 Experimental Setup
Before passing the instances to SCIP and starting the actual branch-and-bound process, we apply simple presolving steps [2, 4]. First, we perform some model cleanup, i.e., we delete already satisfied constraints and convert singleton rows to variable bounds. Further, we apply a limited form of constraint propagation in rational arithmetic in order to reduce the number of infinite bounds, which helps to increase the success rate of safe dual bounding.
Event system in SCIP.
In order to track SCIP’s branch-and-bound process, we register so-called events that inform us every time a leaf is reached. In detail, our implementation uses the following events. Every time SCIP finds a solution, it triggers the event BESTSOLUTION found. A NODEINFEASIBLE event is triggered both when the LP relaxation at this leaf is deemed infeasible, or when a node is pruned. The LP solving status can be accessed to differentiate whether the relaxed LP is infeasible or feasible and eventually pruned. Note that when SCIP creates new nodes it adds them to a queue and sets the local lower bound to the objective value of the LP relaxation of its parent. After a new solution is found, SCIP immediately removes all nodes from the queue whose lower bound is worse than the objective of the newly found solution. Each of these removed leaves triggers a NODEDELETE event. Finally, every integer-feasible node triggers an NODEFEASIBLE event.
@@ 3.2 Evaluation of branch-and-bound decisions in SCIP
Table 1 provides a first overview of the results for the floating-point solver SCIP on both test sets FPEasy and NumDiff, summarizing how many instances were correctly solved and on how many SCIP returns an inexact solution or even an incorrect status. The columns “correct” state the number of instances where no errors occurred as defined in Section 2.2: All solutions are either exactly integer feasible or convertible to such, all infeasible leaves are infeasible in rational arithmetic, and any node pruning is also justified in exact arithmetic. By contrast, the columns “fails” state the number of instances where at least one such error occurred and solving finished within the time limit. The column “primal” reports the number of instances with a solution error (see Section 2.2), the column “dual” reports the number of instances with a bound, gap, or infeasibility error.
Note that for two instances (30:70:4_5:0_95:100 and npmv07) SCIP timed out before reaching a leaf or producing a solution and therefore did not issue any event. Since no wrong decision was made these instances are labeled as correct.
@@
Regardless of the permutation, for 55 of the 57 instances of FPEasy, we could verify that SCIP followed a fully correct solving process. On 49 instances it terminated with an exact optimal solution, while on 6 instances SCIP hit the time limit. Only for two instances, we encountered dual fails. On one instance (vpm2) the results is still correct, on the other instance (dano3_4), the result is slightly suboptimal. This can also be seen in Table 5, where we compare the results of floating-point SCIP to the results of a numerically exact version of SCIP [6, 8, 9], on the instances where SCIP made at least one wrong decisions.
As expected, verifying the results of the NumDiff test set leads to more errors. Only on 25 (original) and 27 (permuted) of the 50 instances, respectively, SCIP followed a fully correct solution process. On each seed, 14 of these instances hit the time limit, so only for 11 respectively 13 instances SCIP terminated with a verifiable exact optimal solution. Still, on both test sets, a notable amount of the floating-point solves exhibits no errors.
For a more detailed analysis, Tables 2, 3 and 4 include all instances labeled as “fail” in Table 1 and those instances where SCIP produced a “correct” result but made incorrect decisions during the solving process. All instances where SCIP made a wrong decision on FPEasy are listed in Table 2. Instances of NumDiff that encountered errors and are solved within the time limit are listed in Table 3. A complete overview NumDiff instances with wrong decisions where SCIP timed out are listed in Table 4. The column “leaves” lists the total number of processed leaves of the tree; the remaining columns report the number of leaves with errors according to the definition in Section 2.2, where “W” and “S” stands for weak and strong errors, respectively.
Sol Bound Gap
@@
Notably, only the instances ns1859355 and vpm2 show a correct result despite wrong decisions during the solving process. On FPEasy the vast majority of instances in Table 2 are solved without wrong decisions. SCIP incorrectly evaluated nodes on only two instances, affecting one node in dano3-4 and 39/178 nodes in vpm2, respectively. Though 39 or 178 wrong decisions on vpm2 seems to be a notable amount of wrong decisions, compared to the overall evaluated nodes on the entire FPEasy test set the wrong decisions on these two instances make up a tiny portion.

### v3

@@ 2 Numerical Errors in LP-based Branch and Bound
where we assume rational input data $A\in\mathbb{Q}^{m\times n}$, $c\in\mathbb{Q}^{n}$, $u,\ell\in(\mathbb{Q}\cup\{\pm\infty\})^{n}$, and $b\in\mathbb{Q}^{m}$. The index set $\mathcal{I}\subseteq\{1,\dots,n\}$ defines which decision variables are required to be integer.
MIPs are typically solved by variants of LP-based branch and bound, which dates back to [16] and still forms the backbone of today’s most competitive MIP solvers. In Section 2.1, we describe briefly the procedure of an LP-based branch-and-bound algorithm. For a more detailed description we refer to [1]. In Section 2.3, we give an overview of the techniques we use to obtain an exact evaluation of a node that was solved in floating-point arithmetic. In Section 2.2, we categorize how numerics can impact the critical branch-and-bound decisions in a floating-point solver and how incorrect decisions affect the overall result.
@@ 2.1 Basics of the LP-based Branch-and-Bound Algorithm
The first step in an LP-based branch-and-bound algorithm is to relax the problem by dropping all integrality constraints. The resulting LP relaxation can be solved by an LP solver. If the LP solution is integer feasible, i.e., all integer variables have an integral solution value, the problem is solved. If the LP relaxation is detected to be infeasible, then also the original MIP is infeasible. Otherwise, the algorithm branches: It chooses an integer variable $x_{i}$ with a fractional value $\hat{x}_{i}$ in the solution of the LP relaxed and creates two new sub-problems by taking the original problem and adding the constraint $x_{i}\geq\lceil\hat{x}_{i}\rceil$ and $x_{i}\leq\lfloor\hat{x}_{i}\rfloor$, respectively. This ensures that the optimal solution is preserved in at least one of the two sub-problems and the current LP solution is no longer feasible in both sub-problems. This process of LP solving and branching is iterated recursively.
If during this process an integer feasible solution is found its objective value can be used to update the primal bound, which is the objective value of the best known solution encountered so far. This primal bound can be used to prune, i.e., remove from further consideration all open sub-problems for which thedual bound, i.e., the objective value of the sub-problem’s LP relaxation, is greater than or equal to the primal bound.
This recursive procedure produces a search tree with sub-problems as nodes. A node is called leaf if it is not further branched on because it is (a) LP infeasible, (b) it was pruned, or (c) the LP solution is integer feasible. Figure 1 presents an example of a branch-and-bound tree for a minimization problem. The solution of the relaxed LP at the node is displayed slightly above and to the right of the node, while its objective value is positioned on the left side. The leaf $N_{4}$ is integer feasible and appears to be also optimal. The leaf $N_{3}$ is LP infeasible and the node $N_{1}$ is cut off since branching on $N_{1}$ does not improve the solution of $N_{4}$ anymore.
@@
[…]
@@ 2.2 Types of Numerical Errors in LP-based Branch and Bound
By design, a floating-point solver may produce approximate solutions with slight violations of integrality and constraint feasibility. A simple post-processing step is to first round the value of all integer variables to the nearest integer and second, if continuous variables are present, to compute rational values for these such that all constraints are satisfied exactly. If this is certifiably not possible, i.e., if the floating-point solver accepted an assignment for the integer variables as feasible that cannot be completed to an exactly feasible solution, we refer to this as a solution error.
@@
If a floating-point solver uses a lower bound to prune a node, this decision can be verified a posteriori by computing solving the LP relaxation exactly or computing a safe dual bound that is valid in exact arithmetic, by techniques described in Section 2.3. If such a safe bound is greater than or equal to the current primal bound, then the decision to prune the node was correct. Otherwise, if the exact objective value of the LP relaxation is below the primal bound, we label this decision as a bound error. Again, we distinguish two cases. If later during the solving process an exactly feasible solution is found that improves the primal bound and in hindsight justifies pruning the leaf, we call the error a weak bound error, otherwise a strong bound error.
@@
For the third type of error, consider the situation that a floating-point solver declares a node to be integer feasible and produces an approximate solution that can be converted to an exactly feasible solution after rounding the integer assignment. Although we record no solution error here, the decision to terminate the search below this node may still be incorrect if the objective value of the converted, exact solution does not match the dual bound derived by solving the LP relaxation exactly. We refer to this error as a gap error, and as above we distinguish between weak gap errors that are justified in hindsight and strong gap errors.
@@
Finally, if the floating-point solver declares a leaf infeasible but there exists a solution to the LP relaxation that is exactly feasible we refer to this situation as an infeasibility error.
@@
Also in the presence of strong bound or strong gap errors, the solution or at least its objective value $z^{*}$ returned by the floating-point solver may still be optimal. Because our analysis produces a safe dual bound at all occurrences of errors outlined above, we can then take the minimum of all these dual bounds and derive a global lower bound $\hat{z}<z^{*}$ for the true optimal objective value. This certifies that the true optimal objective value must lie in the interval $[\hat{z},z^{*}]$. If infeasibility errors occur, i.e., if a leaf is removed although the LP relaxation is feasible in exact arithmetic, we can use the exact objective value of the LP relaxation as a valid dual bound.
@@ 2.3 Checking Numerical Correctness of Floating-point Decisions
A straightforward way to evaluate whether and which of the errors described above occurs at a leaf node is to solve the LP relaxation once again with an exact LP solver and compare the results. For most but small instances, this approach is prohibitively slow. Hence, we use a hierarchy of methods that may be able to verify or refute the correctness of leaf decisions faster, described in the following.
@@
Consider the dual of the LP relaxation,
$\max\big{\{}b^{T}y+\ell^{T}r^{+}-u^{T}r^{-}:A^{T}y+r^{+}-r^{-}=c,\,y,r^{+},r^{-}\geq 0\big{\}},$
and assume we have an approximate dual solution $(\hat{y},\hat{r}^{+},\hat{r}^{-})$ with dual residual error
$\hat{\varepsilon}:=c-A^{T}\hat{y}-\hat{r}^{+}+\hat{r}^{-}\in\mathbb{Q}^{n}.$ (1)
If we compute, in exact arithmetic,
$r^{+}_{i}:=\hat{r}^{+}_{i}+\max\{\hat{\varepsilon}_{i},0\}\textnormal{ and }r^{-}_{i}:=\hat{r}^{-}_{i}-\min\{\hat{\varepsilon}_{i},0\}$ (2)
for all variables $i\in\{1,\ldots,n\}$, then $\{\hat{y},r^{+},r^{-}\}$ is an exactly feasible dual solution with objective value $b^{T}\hat{y}+\ell^{T}r^{+}-u^{T}r^{-}$. This modification thus ensures dual feasibility at the cost of reducing the objective value by $\ell^{T}(\hat{r}^{+}-r^{+})+u^{T}(r^{-}-\hat{r}^{-})$. The quality of the resulting safe dual bound is impacted by the variable bounds $u,\ell$, which, ideally, should be as tight as possible to minimize the amount by which the approximate dual bound is weakened. If infinite variable bounds are used during the computation, the resulting safe bound becomes negative infinity. This technique is first explained in [17] in a version using interval arithmetic and discussed in more detail in [21]. Safe bounding can be applied analogously to validate a Farkas proof for infeasible LPs.
Rational reconstruction.
In the hope of recovering an exact rational solution value from an approximate floating-point value, we use continued fractions approximation, which can be computed by the extended Euclidean algorithm. This way, we can round floating-point values to a nearby rational number with a limited denominator. This method is applied to both primal and dual LP solutions, as well as to Farkas proofs. Originally, this technique is described and improved in [18, 20, 23, 25]. We rely on an implementation in the exact LP solver SoPlex [13, 14].
Factorization.
If a simplex basis is available for the LP relaxation solution, we perform an exact rational LU factorization in order to compute the exact rational solution associated to the basis reached by the floating-point LP solver. Again, we use the implementation in SoPlex [13, 14]. This step is only skipped for LP infeasible nodes since no basis can be obtained from the LP solver.
@@
The last and most expensive option to derive the most accurate lower bound on the objective is to call a round-off-error-free LP solver. The open-source solver SoPlex provides a configuration to perform such an exact rational solve [13, 14]. We refer to this fallback as Exact-SoPlex. It is performed only if all techniques described above are not applicable or fail to decide the status of a leaf.
Exact LP solving is also used to test whether an approximate primal solution can be converted to an exact primal solution by fixing the integer variables to the rounded values of the floating-point solution and solving the LP over the continuous variables in rational arithmetic with Exact-SoPlex. For each primal solution value, it is ensured to lie within its lower and upper bounds.
@@ 3 Computational Study
The guiding questions for the design of our experiments are the following: Assuming the final result produced by a floating-point branch-and-bound solver is correct, can we verify that the also reasoning of the individual decisions of the solver is correct? If not, how many of the critical decisions would require re-evaluation, and which of the errors defined in Section 2.2 occur at which frequency? Furthermore, how effective are the different post-processing techniques outlined in Section 2.3. Before evaluating our experiments in Section 3.2, we first describe the setup and specify the software used in Section 3.1.
@@ 3.1 Experimental Setup
Before passing the instances to SCIP and starting the actual branch-and-bound process, we apply simple presolving steps [2, 4]. First, we perform some model cleanup, i.e., we delete redundant constraints and convert singleton rows to variable bounds. Further, we apply a limited form of constraint propagation in rational arithmetic in order to reduce the number of infinite bounds and help increase the success rate of safe dual bounding.
Event system.
In order to track SCIP’s branch-and-bound process, we register so-called events that inform us every time a leaf is reached. In detail, our implementation uses the following events. Every time SCIP finds a solution, it triggers the event BESTSOLUTION. A NODEINFEASIBLE event is triggered both when the LP relaxation at the current leaf is deemed infeasible, or when a node is pruned. We access the LP solving status to differentiate whether the relaxed LP is infeasible or feasible and eventually pruned. Note that when SCIP creates new nodes it adds them to a queue and sets the local lower bound to the objective value of the LP relaxation of its parent. After a new solution is found, SCIP immediately removes all nodes from the queue whose lower bound is worse than the objective of the newly found solution. Each of these removed leaves triggers a NODEDELETE event. Finally, every integer-feasible node triggers a NODEFEASIBLE event.
@@ 3.2 Evaluation of Branch-and-Bound Decisions in SCIP
Table 1 provides a first overview of the results for the floating-point solver SCIP on both test sets FPEasy and NumDiff, summarizing how many instances were correctly solved and on how many SCIP returns an inexact solution or even an incorrect status. The columns “correct” state the number of instances where no errors occurred as defined in Section 2.2: All solutions are either exactly integer feasible or convertible to such, all infeasible leaves are infeasible in rational arithmetic, and any node pruning is also justified in exact arithmetic. By contrast, the columns “fails” state the number of instances where at least one such error occurred and solving finished within the time limit. The column “primal” reports the number of instances with a solution error (see Section 2.2), the column “dual” reports the number of instances with a bound, gap, or infeasibility error. Note that for the two MIPLIB-instances (30:70:4_5:0_95:100 and npmv07) SCIP timed out before reaching a leaf or producing a solution and therefore did not issue any event. Since no wrong decision was made these instances are labeled as correct.
@@
Regardless of the permutation, for 55 of the 57 instances of FPEasy, we could verify that SCIP followed a fully correct solving process. On 49 instances it terminated with an exact optimal solution, while on 6 instances SCIP hit the time limit. Only for two instances, we encountered dual fails. On one instance (vpm2) the result is still correct, on the other instance (dano3_4), the result is slightly suboptimal. This can also be seen in Table 5, where we compare the results of floating-point SCIP to the results of a numerically exact version of SCIP [6, 8, 9], on the instances where SCIP made at least one wrong decisions.
As expected, verifying the results of the NumDiff test set exhibits more errors. Only on 25 (original) and 27 (permuted) of the 50 instances, respectively, SCIP followed a fully correct solution process. On each seed, 14 of these instances hit the time limit, so only for 11 respectively 13 instances SCIP terminated with a verifiably exact optimal solution. Still, on both test sets, a notable amount of the floating-point solves exhibits no errors.
For a more detailed analysis, Tables 2, 3 and 4 include all instances labeled as “fail” in Table 1 and those instances for which SCIP produced a “correct” result but made incorrect decisions during the solving process. All instances where SCIP made a wrong decision on FPEasy are listed in Table 2. Instances of NumDiff that encountered errors and are solved within the time limit are listed in Table 3. A complete overview NumDiff instances with wrong decisions where SCIP timed out are listed in Table 4. The column “leaves” lists the total number of processed leaves of the tree; the remaining columns report the number of leaves with errors according to the definition in Section 2.2, where “W” and “S” stands for weak and strong errors, respectively.
Sol Bound Gap Inf
@@
Notably, only the instances ns1859355 and vpm2 show a correct result despite wrong decisions during the solving process. On FPEasy the vast majority of instances in Table 2 are solved without wrong decisions. SCIP incorrectly evaluates nodes on only two instances, affecting one node in dano3-4 and 39 of 178 nodes in vpm2, respectively. Though 39 or 178 wrong decisions on vpm2 seems to be a notable amount of wrong decisions, compared to the overall evaluated nodes on the entire FPEasy test set the wrong decisions on these two instances make up a tiny portion.